OpenStudy (anonymous):

The average value of a certain type of automobile was $14,040 in 1994 and depreciated to $4140 in 1999. Let y be the average value of the automobile in the year x, where x = 0 represents 1994. Write a linear equation that models the value of the automobile in terms of the year x. a. y = -1980x + 14,040 b. y = -1980x - 5760 c. y = - 1/1980 x - 4140 d. y = -1980x + 4140

OpenStudy (sriram):

if one of them is correct it has to beA

OpenStudy (anonymous):

If x=0 represents the year 1994, then \(x=5\) represents the year 1999. So, we have the the two points (0,14,040) and (5,4140) in the x-y plane.

OpenStudy (anonymous):

\(y=mx+b\) is the linear equation that represents the model, where m is the slope and b is the y-intercept. b is known to be \(14,040\). m can be found as \(b=\frac{4140-14,040}{5-0}=-1,980\). Hence \(y=-1980x+14040\), which is answer a.